OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? x(x + 8) = 9 28 64 100

OpenStudy (anonymous):

first, multiply out and then move everything to one side. identify a, b and c and then go from there.

OpenStudy (gorv):

x(x+8)=9 x^2+8x-9=0

OpenStudy (gorv):

compare with standard eqn ax^2+bx+c=0

OpenStudy (anonymous):

@venkat409 please read the code of conduct and don't just give out answers!

OpenStudy (gorv):

a=1 b=8 c=-9

OpenStudy (gorv):

so now put these values in your eqn and tell what u got